Find three consecutive integers such that three times the largest increased by two is equal to five times the smallest increased by three times the middle integer.
Answer by mananth(16946)

let n (n+1) & (n+2) be the numbers
3(n+2) +2 = 5(n) +3(n+1)

3n+6 +2 = 5n+3n+3
3n+8 =8n+3
5n =5
n =1
1,2,3 are the numbers